# Truth in the Few: High-Value Data Selection for Efficient Multi-Modal Reasoning

### v1

@@ Abstract
While multi-modal large language models (MLLMs) have made significant progress in complex reasoning tasks via reinforcement learning, it is commonly believed that extensive training data is necessary for improving multi-modal reasoning ability, inevitably leading to data redundancy and substantial computational costs. However, can smaller high-value datasets match or outperform full corpora for multi-modal reasoning in MLLMs? In this work, we challenge this assumption through a key observation: meaningful multi-modal reasoning is triggered by only a sparse subset of training samples, termed cognitive samples, whereas the majority contribute marginally. Building on this insight, we propose a novel data selection paradigm termed Reasoning Activation Potential (RAP), which identifies cognitive samples by estimating each sample’s potential to stimulate genuine multi-modal reasoning by two complementary estimators: 1) Causal Discrepancy Estimator (CDE) based on the potential outcome model principle, eliminates samples that overly rely on language priors by comparing outputs between multi-modal and text-only inputs; 2) Attention Confidence Estimator (ACE), which exploits token-level self-attention to discard samples dominated by irrelevant but over-emphasized tokens in intermediate reasoning stages. Moreover, we introduce a Difficulty-aware Replacement Module (DRM) to substitute trivial instances with cognitively challenging ones, thereby ensuring complexity for robust multi-modal reasoning. Experiments on six datasets show that our RAP method consistently achieves superior performance using only 9.3% of the training data, while reducing computational costs by over 43%. Our code is available at https://github.com/Leo-ssl/RAP.
@@
1 Introduction
Improving complex reasoning in multi-modal large language models (MLLMs) qwen2025qwen25technicalreport ; openai2023gpt4 remains a fundamental challenge. While large-scale reinforcement learning (RL) guo2025deepseek ; hu2024openrlhf ; r1vDBLP:journals/corr/abs-2503-10615 has shown promise in enhancing reasoning capability, the prevailing assumption openai2024reasoning ; kimiteam2025kimik15scalingreinforcement suggests that scaling training data is a necessary condition for developing advanced reasoning ability, thus leading to data redundancy and substantial training costs. Recent studies s1 ; limr indicate that LLMs trained on high-quality curated datasets can outperform those trained on full corpora. However, it remains unclear whether this principle generalizes to multi-modal contexts, where effective cross-modal integration is important. This raises a critical question: can smaller high-value data achieve competitive or superior multi-modal reasoning compared to training on full corpora in MLLMs ? To investigate this, as shown in Figure 1(a), we empirically analyze the effect of data scale on multi-modal reasoning performance. Notably, training with only 20% of the data leads to merely a 0.8% performance degradation compared to the full dataset, suggesting that indiscriminate data scaling may have minimal or even negative effects. We hypothesize that such data augmentation diminishes the influence of high-value samples, termed cognitive samples, which are essential for guiding effective cross-modal integration during reasoning.
@@ 1 Introduction
To validate this assumption, we analyze the characteristics of training samples and find that most fail to encourage joint attention to both modalities during reasoning. Specifically, we identify two main types of ineffective samples: 1) Language-prior biased samples (Figure 2(a)), where the model produces nearly indistinguishable outputs given text-only and multi-modal inputs due to over-reliance on language priors lp1 ; vcd . Such samples enable models to solve tasks with minimal utilization of visual semantics, thus impairing their ability to cross-modal integration. (2) Attention-biased samples (Figure 2(b)), where the model over-attends to semantically irrelevant tokens (e.g., the punctuation “.”), thereby obstructing the exploration of crucial cross-modal relationships. The above findings highlight the need to prioritize cross-modal interactions in data selection for multi-modal reasoning. However, existing methods rely on unimodal textual quality, such as human-annotated difficulty estimation s1 ; limr or reward-based sampling limo . These approaches not only incur substantial manual annotation costs s1 ; limo and considerable filtering time limr , but also fail to estimate whether samples effectively facilitate cross-modal integration.
@@
However, relying solely on output-level measures neglects the reliability of internal reasoning dynamics. Therefore, we propose the Attention Confidence Estimator (ACE) to model the quality of internal reasoning behavior based on token-level attention distributions, thus removing attention-biased samples characterized by high attention to irrelevant tokens. Despite their efficacy, combining these two estimators alone might retain overly simplistic samples while discarding challenging yet valuable instances, thereby constraining the model’s reasoning upper bound. To address this limitation, we propose a Difficulty-aware Replacement Module (DRM) to replace trivial samples with suitable challenging alternatives, which ensure sufficient data complexity for robust multi-modal reasoning. Finally, the results in Figure 1(b) demonstrate that RAP achieves state-of-the-art performance with only 5,159 samples, compared to the full dataset of 54,931 samples, while reducing training costs by over 43%. These findings validate our insight that data quality is more important than blind data scaling for multi-modal reasoning in RL, revealing the “truth in the few” phenomenon.
Our main contribution: 1) We reveal a “truth in the few” phenomenon that smaller high-quality datasets can outperform full corpora for multi-modal reasoning in MLLMs. 2) We propose two novel estimators: a Causal Discrepancy Estimator (CDE) to eliminate samples that overly rely on language priors, and an Attention Confidence Estimator (ACE) to filter out attention-biased samples with irrelevant semantic focus. 3) We introduce a Difficulty-aware Replacement Module (DRM) to preserve sufficient data complexity, effectively improving the model’s reasoning performance ceiling.
@@
2 Related Work
@@
Reinforcement Learning for (M)LLM Reasoning. Reinforcement learning (RL), initially popularized by RL from Human Feedback (RLHF) hu2024openrlhf ; rlhf1 , has become a key paradigm for improving the reasoning ability of both LLMs and MLLMs. Recent methods step-dpo ; PR-DPO ; dpo extend RL beyond human preference alignment, explicitly achieving reasoning improvement via policy-gradient algorithms such as Proximal Policy Optimization (PPO) ppo and reward-centric optimizations including Reinforce Leave-One-Out (RLOO) rloo and Group Relative Policy Optimization (GRPO) guo2025deepseek . Moreover, several methods lmmr1 ; r1vDBLP:journals/corr/abs-2503-10615 ; vrftliu2025visual ; mpowang2024enhancing explore the use of RL to enhance the visual reasoning in MLLMs. However, these methods typically rely on large-scale data mulberry ; MM-Eureka , which fails to adequately consider the quality of training samples. To this end, we introduce a novel RAP method to select valuable samples, ensuring that the training process stimulates effective multi-modal reasoning.
@@
Data Selection for Reasoning. Traditional approaches guo2025deepseek ; kimiteam2025kimik15scalingreinforcement ; mulberry generally highlight the importance of data scaling, suggesting that larger data volumes can lead to better model performance. Contrary to this, recent methods limo ; limr ; s1 demonstrate that curated datasets can outperform those trained on full corpora in complex textual reasoning tasks. For example, LIMO limo and s1 s1 demonstrate that models trained with just 817 or 1000 curated samples achieve better performance compared to models trained on larger datasets. Inspired by these, we focus on whether and how a minimal but valuable dataset can similarly enhance multi-modal reasoning within multi-modal contexts.
Causal Mechanism. Recent advances have demonstrated the significance of causal mechanisms in complex reasoning tasks casualnipsDBLP:conf/nips/RohekarGN23 ; casualrw2DBLP:conf/aaai/ZhangZZ24 . To better capture causal relationships in the reasoning processes of MLLMs, prior studies have used existing causal inference frameworks, such as Structural Causal Models (SCM) scholkopf2012causal and Potential Outcome Models (POM) pom . For example, the POM offers a rigorous analytical structure to examine causal dependencies casualnipsDBLP:conf/nips/RohekarGN23 ; pom-llm , facilitating the isolation of causal factors directly influencing model reasoning. Therefore, we adopt the POM framework to systematically quantify differences in MLLM outputs given multi-modal and text-only inputs.
3 Method
@@
Overview of RAP Method. As shown in Figure 3, our RAP method aims to identify high-value training samples $x_{cd}$, termed cognitive samples, that effectively stimulate multi-modal reasoning in MLLMs during RL training. Given a training instance $x=(x_{t},x_{v})$, we estimate its reasoning activation potential from two perspectives. First, we adopt the potential outcome model to quantify the output-level discrepancy $D$ between the model predictions under multi-modal inputs, $\textit{{Y}}_{1}(x)$, and text-only inputs, $\textit{{Y}}_{0}(x_{t})$. Samples with low discrepancy values be considered as language-prior biased and are discarded accordingly. Second, we compute a confidence score $\psi(\textit{{A}})$ from the self-attention matrix $\textit{{A}}\in\mathbb{R}^{d\times d}$ of the model’s final layer to assess the model’s focus on semantically meaningful tokens. Samples with attention focused on irrelevant tokens, below a threshold $\lambda_{a}$, are excluded as attention-biased. Moreover, we replace overly easy samples with an equal number of hard examples potentially missed by the initial model due to limited reasoning capacity, thereby enhancing the model’s reasoning upper bound. The resulting cognitive samples form a refined training dataset that supports more efficient and robust multi-modal reasoning in MLLMs. Finally, we utilize these cognitive samples $x_{cd}$ to optimize the model by maximizing the objective of GRPO guo2025deepseek ; hintgrpo .
3.1 Causal Discrepancy Estimator
@@
Background of Potential Outcome Model. The foundation of causal inference is rooted in the Neyman-Rubin POM rubin1974estimating ; pom , which aims to estimate the effect of a treatment $T$ on an outcome Y for individuals described by covariates X. In this work, we consider the treatment variable $T$ to be binary, i.e., $T\in\{0,1\}$. Under this framework, each unit $u$ is associated with two potential outcomes: $\textit{{Y}}_{1}(u)$, the outcome if the unit receives the treatment ($T=1$), and $\textit{{Y}}_{0}(u)$, the outcome under control ($T=0$). The Individual Treatment Effect (ITE) ite1 ; ite2 is defined as the difference $\textit{{Y}}_{1}(u)-\textit{{Y}}_{0}(u)$. However, due to the fundamental problem of causal inference itedrawpearl2009causal ; itedraw2hammerton2021causal , only one of these outcomes can ever be observed for a given individual, rendering the ITE fundamentally unidentifiable. To address this, the prior work CATEabrevaya2015estimating proposes the Conditional Average Treatment Effect (CATE), which represents the expected treatment effect conditioned on covariates:
@@
$D(x)=\mathbb{E}\left[\mathbb{I}(\textit{{Y}}_{1}=\textit{{Y}}_{g})-\mathbb{I}(\textit{{Y}}_{0}=\textit{{Y}}_{g})\mid x\right]=\frac{1}{M}\sum_{i=1}^{M}\left[\mathbb{I}(Y_{1}(x^{(i)})=Y_{g}^{(i)})-\mathbb{I}(Y_{0}(x_{t}^{(i)})=Y_{g}^{(i)})\right],$ (4)
@@
3.2 Attention Confidence Estimator
While the CDE effectively identifies samples requiring multi-modal reasoning from an output-level perspective, it does not assess the quality of internal reasoning processes. Recent studies opera24 ; anctokDBLP:conf/emnlp/WangLDCZMZS23 reveal an insightful phenomenon: tokens that receive excessive attention weights can dominate the prediction without using a meaningful semantic context. Motivated by this, we explicitly quantify the internal reasoning quality via self-attention distributions, thus filtering out attention-biased samples.

### v2

@@ Abstract
While multi-modal large language models (MLLMs) have made significant progress in complex reasoning tasks via reinforcement learning, it is commonly believed that extensive training data is necessary for improving multi-modal reasoning ability, inevitably leading to data redundancy and substantial computational costs. However, can smaller high-value datasets match or outperform full corpora for multi-modal reasoning in MLLMs? In this work, we challenge this assumption through a key observation: meaningful multi-modal reasoning is triggered by only a sparse subset of training samples, termed cognitive samples, whereas the majority contribute marginally. Building on this insight, we propose a novel data selection paradigm termed Reasoning Activation Potential (RAP), which identifies cognitive samples by estimating each sample’s potential to stimulate genuine multi-modal reasoning by two complementary estimators: 1) Causal Discrepancy Estimator (CDE) based on the potential outcome model principle, eliminates samples that overly rely on language priors by comparing outputs between multi-modal and text-only inputs; 2) Attention Confidence Estimator (ACE), which exploits token-level self-attention to discard samples dominated by irrelevant but over-emphasized tokens in intermediate reasoning stages. Moreover, we introduce a Difficulty-aware Replacement Module (DRM) to substitute trivial instances with cognitively challenging ones, thereby ensuring complexity for robust multi-modal reasoning. Experiments on six datasets show that our RAP method consistently achieves superior performance using only 9.3% of the training data, while reducing computational costs by over 43%.
@@
I Introduction
Improving complex reasoning in multi-modal large language models (MLLMs) [1, 2] remains a fundamental challenge. While large-scale reinforcement learning (RL) [3, 4, 5] has shown promise in enhancing reasoning capability, the prevailing assumption [6, 7] suggests that scaling training data is a necessary condition for developing advanced reasoning ability, thus leading to data redundancy and substantial training costs. Recent studies [8, 9] indicate that LLMs trained on high-quality curated datasets can outperform those trained on full corpora. However, it remains unclear whether this principle generalizes to multi-modal contexts, where effective cross-modal integration is important. This raises a critical question: can smaller high-value data achieve competitive or superior multi-modal reasoning compared to training on full corpora in MLLMs? To investigate this, as shown in Fig. 1(a), we empirically analyze the effect of data scale on multi-modal reasoning performance. Notably, training with only 20% of the data leads to merely a 0.8% performance degradation compared to the full dataset, suggesting that indiscriminate data scaling may have minimal or even negative effects. We hypothesize that such data augmentation diminishes the influence of high-value samples, termed cognitive samples, which are essential for guiding effective cross-modal integration during reasoning.
@@ I Introduction
To validate this assumption, we analyze the characteristics of training samples and find that most fail to encourage joint attention to both modalities during reasoning. Specifically, we identify two main types of ineffective samples: 1) Language-prior biased samples (Fig. 2(a)), where the model produces nearly indistinguishable outputs given text-only and multi-modal inputs due to over-reliance on language priors [10, 11]. Such samples enable models to solve tasks with minimal utilization of visual semantics, thus impairing their ability to cross-modal integration. (2) Attention-biased samples (Fig. 2(b)), where the model over-attends to semantically irrelevant tokens (e.g., the punctuation “.”), thereby obstructing the exploration of crucial cross-modal relationships. The above findings highlight the need to prioritize cross-modal interactions in data selection for multi-modal reasoning. However, existing methods rely on unimodal textual quality, such as human-annotated difficulty estimation [8, 9] or reward-based sampling [12]. These approaches not only incur substantial manual annotation costs [8, 12] and considerable filtering time [9], but also fail to estimate whether samples effectively facilitate cross-modal integration.
@@
However, relying solely on output-level measures neglects the reliability of internal reasoning dynamics. Therefore, we propose the Attention Confidence Estimator (ACE) to model the quality of internal reasoning behavior based on token-level attention distributions, thus removing attention-biased samples characterized by high attention to irrelevant tokens. Despite their efficacy, combining these two estimators alone may retain overly simplistic samples while discarding challenging yet valuable instances, thereby constraining the model’s reasoning upper bound. To address this limitation, we propose a Difficulty-aware Replacement Module (DRM) to replace trivial samples with suitable challenging alternatives, which ensure sufficient data complexity for robust multi-modal reasoning. Finally, the results in Fig. 1(b) demonstrate that our RAP method achieves state-of-the-art performance with only 5,159 samples, compared to the full dataset of 54,931 samples, while reducing training costs by over 43%. These findings validate our insight that data quality is more important than blind data scaling for multi-modal reasoning in RL, revealing the “truth in the few” phenomenon.
Overall, our contributions can be summarized as follows:
  • We reveal a “truth in the few” phenomenon that smaller high-quality datasets can outperform full corpora for multi-modal reasoning in LVLMs.
  • We propose two novel estimators: a Causal Discrepancy Estimator (CDE) to eliminate samples that overly rely on language priors, and an Attention Confidence Estimator (ACE) to filter out attention-biased samples with irrelevant semantic focus.
  • We introduce a Difficulty-aware Replacement Module (DRM) to preserve sufficient data complexity, effectively improving the model’s reasoning performance ceiling.
@@
II Related Work
II-A Reinforcement Learning for (M)LLM Reasoning
@@
Reinforcement learning (RL), initially popularized by RL from Human Feedback (RLHF) [4, 13], has become a key paradigm for improving the reasoning ability of both LLMs [14] and MLLMs [15, 16]. Recent methods [17, 18, 19] extend RL beyond human preference alignment, explicitly achieving reasoning improvement via policy-gradient algorithms such as Proximal Policy Optimization (PPO) [20] and reward-centric optimizations including Reinforce Leave-One-Out (RLOO) [21] and Group Relative Policy Optimization (GRPO) [3]. Moreover, several methods [22, 5, 23, 24] explore the use of RL to enhance the visual reasoning in MLLMs. Recent works have extended RL beyond alignment for enhancing reasoning capability, by utilizing policy gradient algorithms PPO [20] and optimizing reward signals, such as RLOO [21] and GRPO [3]. However, these methods typically rely on large-scale data [25, 26], which fails to adequately consider the quality of training samples. To this end, we introduce a novel RAP method to select valuable samples, ensuring that the training process stimulates effective multi-modal reasoning.
II-B Data Selection for Reasoning
@@
Traditional approaches [3, 7, 25, 27] generally highlight the importance of data scaling, suggesting that larger data volumes can lead to better model performance. In contrast, recent methods [12, 9, 8] demonstrate that carefully curated datasets can outperform models trained on full corpora for challenging textual reasoning tasks. For example, LIMO [12] and s1 [8] show that models trained with only 817 or 1000 curated samples can achieve better performance than those trained with substantially larger datasets, indicating that sample quality and representativeness may be more decisive than sheer quantity. Similarly, LIMR [9] suggests that selective data curation during RL training can further improve outcomes with fewer samples by prioritizing instances that provide clearer optimization signals. Inspired by these, we focus on whether and how a minimal but valuable dataset can similarly enhance multi-modal reasoning within multi-modal contexts.
II-C Causal Mechanism
Recent advances have highlighted the role of causal mechanisms in understanding and improving complex reasoning [28, 29, 30, 31]. To capture causal relationships in MLLM reasoning, prior studies have leveraged established causal inference frameworks, such as Structural Causal Models (SCM) [32] and Potential Outcome Models (POM) [33]. Compared with purely correlational analyses, these frameworks facilitate intervention-based comparisons that help isolate factors genuinely responsible for a model’s decisions. In particular, POM provides a rigorous analytical structure to examine causal dependencies [28, 34], enabling the quantification of how an explicit change in input conditions affects model outputs. Therefore, we adopt the POM framework to systematically measure differences in MLLM predictions under multi-modal inputs versus text-only inputs, thereby assessing the causal contribution of visual information to the reasoning process.
III Method
@@
Overview of RAP Method. As shown in Fig. 3, our RAP method aims to identify high-value training samples $x_{cd}$, termed cognitive samples, that effectively stimulate multi-modal reasoning in MLLMs during RL training. Given a training instance $x=(x_{t},x_{v})$, we estimate its reasoning activation potential from two perspectives. First, we adopt the potential outcome model to quantify the output-level discrepancy $D$ between the model predictions under multi-modal inputs, $\textit{{Y}}_{1}(x)$, and text-only inputs, $\textit{{Y}}_{0}(x_{t})$. Samples with low discrepancy values be considered as language-prior biased and are discarded accordingly. Second, we compute a confidence score $\psi(\textit{{A}})$ from the self-attention matrix $\textit{{A}}\in\mathbb{R}^{d\times d}$ of the model’s final layer to assess the model’s focus on semantically meaningful tokens. Samples with attention focused on irrelevant tokens, below a threshold $\lambda_{a}$, are excluded as attention-biased. Moreover, we replace overly easy samples with an equal number of hard examples potentially missed by the initial model due to limited reasoning capacity, thereby enhancing the model’s reasoning upper bound. The resulting cognitive samples form a refined training dataset that supports more efficient and robust multi-modal reasoning in MLLMs. Finally, we utilize these cognitive samples $x_{cd}$ to optimize the model by maximizing the objective of GRPO [3, 35].
Algorithm 1 The Process of Reasoning Activation Potential (RAP) Method.
1:Original dataset $D=\{(x_{t}^{i},x_{v}^{i},y^{i})\}_{i=1}^{N}$, pretrained model $\theta$, output discrepancy threshold $\lambda_{c}$, attention confidence threshold $\lambda_{a}$, training steps $T$
2:Curated cognitive samples $x_{cd}$
3:Stage 1: Causal Discrepancy Estimator (CDE)
4:for all $(x_{t},x_{v},y)\in D$ do
5:  Generate $M$ outputs with multimodal input $(x_{t},x_{v})$ and text-only input $(x_{t})$
6:  Compute output consistency:
7:    $D(x)=\frac{1}{M}\sum_{j=1}^{M}\Big[\mathbb{I}\!\big(Y_{1}^{(j)}=y\big)-\mathbb{I}\!\big(Y_{0}^{(j)}=y\big)\Big]$
8:end for
9:Form discrepancy-filtered set:
10:  $X_{\mathrm{cde}}=\{\,x\mid D(x)\geq\mu_{D}+\lambda_{c}\sigma_{D}\,\}$
11:Stage 2: Attention Confidence Estimator (ACE)
12:Initialize $X_{\mathrm{ace}}\leftarrow\emptyset$
13:for all $x\in X_{\mathrm{cde}}$ do
14:  Run a forward pass and extract final-layer self-attention $A\in\mathbb{R}^{L\times L}$
15:  Compute attention confidence:
16:    $\psi_{j}(A)=\prod_{i=j}^{L}(\sigma\cdot A_{i,j}),\quad\psi(A)=\max_{j}\psi_{j}(A)$
17:  if $\psi(A)\leq\lambda_{a}$ then
18:   $X_{\mathrm{ace}}\leftarrow X_{\mathrm{ace}}\cup\{x\}$
19:  end if
20:end for
21:Stage 3: Difficulty-aware Replacement Module (DRM)
22:for all $x\in X_{\mathrm{ace}}$ do
23:  Compute difficulty:
24:    $D_{\mathrm{diff}}(x)=1-\frac{1}{M}\sum_{j=1}^{M}\mathbb{I}\!\big(\text{output}^{(j)}\text{ is correct}\big)$
25:end for
26:Identify trivial samples:
27:  $X_{\mathrm{easy}}=\{\,x\mid D_{\mathrm{diff}}(x)=0\,\}$
28:Select top $|X_{\mathrm{easy}}|$ challenging samples from $D\setminus X_{\mathrm{ace}}$ by descending difficulty to form $X_{\mathrm{hard}}$
29:Construct final cognitive dataset:
30:  $x_{cd}=(X_{\mathrm{ace}}\setminus X_{\mathrm{easy}})\cup X_{\mathrm{hard}}$
31:Stage 4: Cognitive Dataset Training
32:Train $\theta$ on $x_{cd}$ via reinforcement learning algorithm (GRPO) for $T$ steps
III-A Causal Discrepancy Estimator
@@
Background of Potential Outcome Model. The foundation of causal inference is rooted in the Neyman-Rubin POM [36, 33], which aims to estimate the effect of a treatment $T$ on an outcome Y for individuals described by covariates X. In this work, we consider the treatment variable $T$ to be binary, i.e., $T\in\{0,1\}$. Under this framework, each unit $u$ is associated with two potential outcomes: $\textit{{Y}}_{1}(u)$, the outcome if the unit receives the treatment ($T=1$), and $\textit{{Y}}_{0}(u)$, the outcome under control ($T=0$). The Individual Treatment Effect (ITE) [37, 38] is defined as the difference $\textit{{Y}}_{1}(u)-\textit{{Y}}_{0}(u)$. However, due to the fundamental problem of causal inference [39, 40], only one of these outcomes can ever be observed for a given individual, rendering the ITE fundamentally unidentifiable. To address this, the prior work [41] proposes the Conditional Average Treatment Effect (CATE), which represents the expected treatment effect conditioned on covariates:
@@
$D(x)$ $=\mathbb{E}\!\left[\mathbb{I}(\textit{{Y}}_{1}=\textit{{Y}}_{g})-\mathbb{I}(\textit{{Y}}_{0}=\textit{{Y}}_{g})\mid x\right]$ (4)
$=\frac{1}{M}\sum_{i=1}^{M}\left[\mathbb{I}\!\big(Y_{1}(x^{(i)})=Y_{g}^{(i)}\big)-\mathbb{I}\!\big(Y_{0}(x_{t}^{(i)})=Y_{g}^{(i)}\big)\right],$
@@
III-B Attention Confidence Estimator
While the CDE effectively identifies samples requiring multi-modal reasoning from an output-level perspective, it does not assess the quality of internal reasoning processes. Recent studies [42, 43] reveal an insightful phenomenon: tokens that receive excessive attention weights can dominate the prediction without using a meaningful semantic context. Motivated by this, we explicitly quantify the internal reasoning quality via self-attention distributions, thus filtering out attention-biased samples.